# AffectGPT-R1: Leveraging Reinforcement Learning for Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ 6 Conclusion
This paper proposes AffectGPT-R1, the first work applying reinforcement learning to OV-MER. We introduce various rewards and incorporate length penalties to mitigate reward hacking. Through extensive experiments, we reveal the impact of different rewards and identify the optimal reward design. AffectGPT-R1 demonstrates strong effectiveness in OV-MER, and this framework also achieves state-of-the-art performance on MER-UniBench, a benchmark for generalized emotion understanding. Furthermore, we examine the influence of various modules and hyperparameters, providing valuable insights for future research. By proving the potential of reinforcement learning as a promising approach for emotion understanding, our work will advance the development of this field toward a reinforcement learning era.
[…]
Supplementary Material
A EW-based Metric
@@
In this section, we present a detailed calculation formula for EW-based metrics. The computation method primarily consists of two components: eliminating the influence of synonyms and defining set-level metrics.
A.1 Handling Synonyms
@@
(L1) We map different forms of words to their base form. For example, happier and happiness are mapped to happy. This function is denoted as $F_{l_{1}}(\cdot)$.
(L2) We map synonyms to a unified word. For example, happy and joyful are mapped to happy. This mapping function is denoted as $F_{l_{2}}(\cdot)$.
(L3) Emotion wheel provides natural grouping information [plutchik1980general]. Since there is no consensus on the emotion wheel, we adopt $K=5$ emotion wheels, consistent with prior work [lian2025ov]. Appendix B visualizes all emotion wheels. For each wheel $w_{k}$, we map all outer labels to their corresponding inner labels. This function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
The clustering functions can be summarized as:
@@
A.2 Set-level Metric
Since we do not restrict the number of predicted and annotated emotions, we employ set-level evaluation metrics. Specifically, suppose the dataset contains $N$ samples. For a sample $x_{i}$, the true labels are $\mathbf{Y}_{i}=\{y_{i}^{j}\}_{j=1}^{n_{i}}$ and the predicted labels are $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$. The evaluation metrics are defined as follows:
@@ A.2 Set-level Metric
B Emotion Wheel
The emotion wheel is a structured visual tool designed to categorize and identify human emotions. It organizes emotions into core categories (the inner sections of the wheel) and their nuanced variations (the outer sections), providing intuitive insights into emotional structure. This paper uses five emotion wheels to compute metrics for OV-MER, aligning with prior work [lian2025affectgpt] to ensure fair comparisons.
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5
Figure 3: Emotion wheels. This paper uses five emotion wheels to calculate EW-based metrics.
C Baseline Details
SECap [xu2024secap].
@@ C Baseline Details
SALMONN [tang2023salmonn].
This model is capable of processing both speech and audio. Specifically, it employs Whisper [radford2023robust] to encode speech and BEATs [chen2023beats] to encode audio. The resulting features are then integrated into the LLM via a window-level Q-Former. Additionally, a LoRA adapter is applied to enhance the LLM’s instruction-following and question-answering capabilities.
Qwen-Audio [chu2023qwen].
@@
Otter [li2025otter].
This model is built upon the OpenFlamingo framework [awadalla2023openflamingo] and trained on the MIMIC-IT dataset [li2025otter], demonstrating strong capabilities in multimodal perception, reasoning, and in-context learning.
Chat-UniVi [jin2024chat].
@@
Video-ChatGPT [maaz2024video].
@@
PandaGPT [su2023pandagpt].
It uses ImageBind [girdhar2023imagebind] to map multiple modalities into a shared embedding space, then employs a projector to transform these features into the LLM’s input space, and finally leverages the LLM to generate responses.
OneLLM [han2024onellm].
@@
Emotion-LLaMA [cheng2024emotion].
@@
R1-Omni [zhao2025r1].
It focuses on basic emotion recognition and uses reinforcement learning with verifiable rewards to address this task. Unlike R1-Omni, AffectGPT-R1 shifts the task to open-vocabulary emotion detection. Due to this task difference, we adopt distinct reward functions and face unique challenges, such as reward hacking, that are not present in R1-Omni. It is important to note that R1-Omni is only available as a preprint on arXiv. Following CVPR policy, we are not required to compare with it. Nevertheless, to thoroughly validate the effectiveness of AffectGPT-R1, we include a comparison with R1-Omni, and the experimental results demonstrate our model’s advantages in open-vocabulary emotion detection and generalized emotion understanding. Additionally, unlike R1-Omni, which only releases its inference code, we provide both training and inference code to enhance reproducibility.
AffectGPT [lian2025affectgpt].
@@
D MER-UniBench Details
MER-UniBench is a unified benchmark that covers three typical tasks, including sentiment analysis, basic emotion recognition, and fine-grained emotion detection, with the aim of evaluating the generalized emotion understanding capabilities of different models. In Table 6, we summarize the data components for each task. Please refer to the original paper for more details [lian2025affectgpt].
Table 6: MER-UniBench. This benchmark encompasses three tasks, and this table presents the data components for each task. In total, this benchmark includes 12,799 samples for testing.
Raw Dataset Selected Subset # Samples
@@ D MER-UniBench Details
CMU-MOSI [zadeh2017tensor] Test 686
CMU-MOSEI [zadeh2018multimodal] Test 4,659
CH-SIMS [yu2020ch] Test 457
CH-SIMS v2 [liu2022make] Test 1,034
@@
MER2023 [lian2023mer] MER-MULTI 411
MER2024 [lian2024mer] MER-SEMI 1,169
IEMOCAP [busso2008iemocap] Session5 1,241
MELD [poria2019meld] Test 2,610
@@
OV-MERD+ [lian2025affectgpt] All 532
E Impact Statement
E.1 Social Impact
Emotion plays an important role in mental health, human-computer interaction, and embodied AI. This paper aims to advance the development of MER techniques, thereby promoting their applications and improving our daily lives.
E.2 Ethics Statement
This paper does not collect or annotate new data but utilizes existing datasets to verify the effectiveness of reinforcement learning in OV-MER. The use of all datasets has been authorized by their owners under corresponding usage requirements. Therefore, no ethical issues are raised in this paper.
F Reproducibility Statement
In this paper, we have made every effort to ensure the reproducibility of our work. In the supplementary material, we provide our source code and command lines for cold-start training, reinforcement learning, and inference. Additionally, we include a detailed README file. To comply with CVPR’s anonymity requirements, we have removed all external links and owner names. After acceptance, we will release the full code on GitHub to facilitate further research.

### v3

@@ 6 Conclusion
This paper proposes AffectGPT-R1, a pioneering work that applies reinforcement learning to OV-MER. We introduce various rewards and incorporate length penalties to mitigate reward hacking. Through extensive experiments, we reveal the impact of different rewards and identify the optimal reward design. AffectGPT-R1 shows strong effectiveness in OV-MER and also achieves state-of-the-art performance on MER-UniBench, a benchmark designed for generalized emotion understanding. Furthermore, we investigate the effects of various modules and hyperparameters, offering valuable insights for future research. By demonstrating the potential of reinforcement learning as a powerful paradigm for emotion understanding, our work will advance the development of this field toward a reinforcement learning era.
Impact Statement
Social Impact.
Emotion plays an important role in mental health, human-computer interaction, and embodied AI. This paper aims to advance the development of MER techniques, thereby promoting their applications.
Ethics Statement.
We do not collect or annotate new data, but use existing datasets to verify the effectiveness of reinforcement learning in OV-MER. The use of all datasets has been authorized by their owners under corresponding usage requirements. Therefore, no ethical issues are raised.
Appendix A Reproducibility Statement
In this paper, we have made every effort to ensure the reproducibility of our work. In the supplementary material, we provide our source code and command lines for cold-start training, reinforcement learning, and inference. Additionally, we include a detailed README file. To comply with ICML’s anonymity requirements, we have removed all external links and owner names. After acceptance, we will release the full code on GitHub to facilitate further research.
Appendix B Novelty Statement
In this paper, we present pioneering work that reveals the potential of reinforcement learning in OV-MER. There are currently two main types of emotion recognition tasks: discriminative emotions and generative emotions. The former primarily involves tasks with a constrained label space, such as basic or dimensional emotions (Wang et al., 2023; Sun et al., 2024; Fang et al., 2025). In contrast, the latter imposes no restrictions on the label space, giving rise to the emerging area of open-vocabulary emotions (Lian et al., 2025c). Recent research has shown a growing interest in shifting from discriminative emotions to generative emotions, driven by the need for more nuanced emotional representations (Lian et al., 2025a; Cheng et al., 2024; Zhang et al., 2025a). However, effectively addressing generative emotions remains an open challenge. In this work, we focus on open-vocabulary emotions, a specific subtask within generative emotions, and propose leveraging reinforcement learning as a solution approach. During our investigation, several practical issues emerged, including the design of reward functions and strategies to mitigate reward hacking. By tackling these challenges, our experimental results demonstrate the effectiveness of reinforcement learning in handling generative emotions. Therefore, this paper should not be viewed merely as an application of reinforcement learning, but rather as a significant milestone in advancing generative emotions research. This work will inspire more researchers to consider reinforcement learning as a foundational tool for addressing generative emotions.
Meanwhile, we would like to emphasize that this work differs from R1-Omni (Zhao et al., 2025), another RL-based framework in affective computing. Specifically, R1-Omni focuses on discriminative emotions, whereas AffectGPT-R1 shifts the task to generative emotions. Due to this task difference, we adopt distinct reward functions and face unique challenges, such as reward hacking, that are not present in R1-Omni. Meanwhile, it is important to note that R1-Omni is only available as a preprint on arXiv. Following the ICML policy, we are not required to compare with it. Nevertheless, to thoroughly validate the effectiveness of AffectGPT-R1, we include a comparison with R1-Omni, and the experimental results demonstrate our model’s advantages in open-vocabulary emotion detection and generalized emotion understanding.
Appendix C Emotion Wheel (EW)
The emotion wheel is a structured visual tool designed to categorize and identify human emotions. It organizes emotions into core categories (the inner sections of the wheel) and their nuanced variations (the outer sections), providing intuitive insights into emotional structure. This paper uses five emotion wheels to compute EW-based metrics, aligning with prior work (Lian et al., 2025a) to ensure fair comparisons. Figure 3 illustrates these emotion wheels.
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5
Figure 3: Emotion wheels. This paper uses five emotion wheels to calculate EW-based metrics, ensuring a fair comparison with previous work (Lian et al., 2025a, c).
Appendix D EW-based Metric
@@
In this section, we present a detailed calculation formula for EW-based metrics (Lian et al., 2025c). The computation method primarily consists of two components: eliminating the influence of synonyms and defining set-level metrics.
D.1 Handling Synonyms
@@
(L1) We map different forms of words to their base form. For example, happier and happiness are mapped to happy. To achieve this, for a base word $w$, we first use GPT to generate various forms of $w$, with the output denoted as $l_{1}(w)$. Since different results may be produced across multiple calls, we repeat this process ten times and merge all the generated words, resulting in $\hat{l}_{1}(w)$. To map these different word forms back to their base form, we simply map $\hat{l}_{1}(w)$ to $w$, and we refer to this mapping function as $F_{l_{1}}(\cdot)$.
(L2) We map synonyms to a unified word. For example, happy and joyful are mapped to happy. To achieve this, similar to L1, we use GPT to generate synonyms for a given word $w$, repeat the process multiple times, and then merge all the generated words, obtaining $\hat{l}_{2}(w)$. To map the synonyms to their unified form, we map each element in $\hat{l}_{2}(w)$ to $w$, and we refer to this mapping function as $F_{l_{2}}(\cdot)$.
(L3) The emotion wheel offers natural grouping information for emotions (Plutchik, 1980). Since there is no consensus on the emotion wheel, we adopt $K=5$ distinct emotion wheels, consistent with prior work (Lian et al., 2025a, c) to ensure fair comparisons. Figure 3 visualizes all five emotion wheels. For each wheel $w_{k}$, we map all outer labels to their corresponding inner labels. This function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
Finally, the clustering functions can be summarized as:
@@
D.2 Set-level Metric
Since we do not restrict the number of predicted and annotated emotions, we employ set-level evaluation metrics. Specifically, suppose the dataset contains $N$ samples. For a sample $x_{i}$, the true labels are $\mathbf{Y}_{i}=\{y_{i}^{j}\}_{j=1}^{n_{i}}$ and the predicted labels are $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$. Here, $n_{i}$ and $\hat{n}_{i}$ are the number of predicted and annotated emotions, respectively. The evaluation metrics are defined as follows:
@@ D.2 Set-level Metric
Appendix E Details of Alignment Reward
E.1 Emotion Word Extraction
To compute the alignment reward, we need to extract emotion words $e^{t}$ from $o^{t}$. To do so, we input $o^{t}$ into Qwen2.5-7B using the following prompt: Please assume the role of an expert in the field of emotions. We provide clues related to the emotional state of a character. Based on the provided clues, please identify the character’s emotional states. Separate different emotional categories with commas, and output only the clearly identifiable emotional categories in a list format. If no emotions can be identified, please return an empty list.
E.2 Similarity Calculation
Alignment reward involves computing the similarity between two emotion sets. In practice, we use the Jaccard Similarity Coefficient, which measures the similarity between two sets by comparing the size of their intersection to the size of their union. For clarity, we adopt the notations defined in Appendix D, and the method for calculating the Jaccard Similarity Coefficient is provided below:
$\mbox{Similarity}_{\mbox{s}}^{k}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|G_{w_{k}}(\mathbf{Y}_{i})\cap G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|}{\left|G_{w_{k}}(\mathbf{Y}_{i})\cup G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|},$ (21)
$\mathrm{is\_similar}\left(\mathbf{Y}_{i},\mathbf{\hat{Y}}_{i}\right)=\frac{1}{K}\sum_{k=1}^{K}\mbox{Similarity}_{\mbox{s}}^{k}.$ (22)
Appendix F Baseline Details
SECap (Xu et al., 2024).
@@ Appendix F Baseline Details
SALMONN (Tang et al., 2023).
This model is capable of processing both speech and audio. Specifically, it employs Whisper (Radford et al., 2023) to encode speech and BEATs (Chen et al., 2023b) to encode audio. The resulting features are then integrated into the LLM via a window-level Q-Former. Additionally, a LoRA adapter is applied to enhance the LLM’s instruction-following and question-answering capabilities.
Qwen-Audio (Chu et al., 2023).
@@
Otter (Li et al., 2025a).
This model is built upon the OpenFlamingo framework (Awadalla et al., 2023) and trained on the MIMIC-IT dataset (Li et al., 2025a), demonstrating strong capabilities in multimodal perception, reasoning, and in-context learning.
Chat-UniVi (Jin et al., 2024).
@@
Video-ChatGPT (Maaz et al., 2024).
@@
PandaGPT (Su et al., 2023).
It uses ImageBind (Girdhar et al., 2023) to map multiple modalities into a shared embedding space, then employs a projector to transform these features into the LLM’s input space, and finally leverages the LLM to generate responses.
OneLLM (Han et al., 2024).
@@
R1-Omni (Zhao et al., 2025).
This framework also leverages reinforcement learning for emotion recognition. Unlike AffectGPT-R1, which focuses on open-vocabulary emotions, this work targets basic emotions. Due to the different objectives, the reward functions used in R1-Omni differ from those employed in AffectGPT-R1. A more detailed comparison between R1-Omni and AffectGPT-R1 is provided in Appendix B.
Emotion-LLaMA (Cheng et al., 2024).
@@
AffectGPT (Lian et al., 2025a).
@@
Appendix G Dataset Details
G.1 Training Datasets
MER-Caption+ (Lian et al., 2025a).
MER-Caption+ is a large-scale descriptive emotion dataset comprising approximately 31K samples and 2K emotion labels. To reduce annotation costs, it introduces a novel model-led human-assisted strategy that leverages human priors to guide both description generation and sample filtering. Specifically, in the description generation stage, the pre-trained models are carefully selected to maximize performance on fine-grained emotion detection. During sample filtering, a two-level filtering process is employed to remove samples with mismatched audio and video, abnormally long descriptions, and those that perform poorly on sentiment analysis and basic emotion recognition.
MER2025-OV (Lian et al., 2025b).
MER2025-OV is specifically designed for OV-MER. Unlike the automatically annotated MER-Caption+, MER2025-OV is a high-quality dataset created through a purely manual annotation process. It employs a multi-round annotation strategy. Initially, a pool of candidate emotions is provided, and four experts in affective computing are asked to select the correct labels and add any missing ones. All selected labels are combined to form the candidate set for the next round. In the second round, four annotators perform another round of labeling, and only labels selected by at least two annotators are retained. This two-stage approach ensures both completeness, by accepting labels chosen by any annotator in the first round, and accuracy, by filtering for consensus in the second. Through this rigorous process, MER2025-OV guarantees the high quality of its annotated emotion labels.
G.2 Testing Datasets
OV-MERD+ (Lian et al., 2025a).
OV-MERD+ is a dataset specifically designed for OV-MER, which is an extended version of OV-MERD (Lian et al., 2025c). The original OV-MERD dataset consists of 332 samples, all of which are fully annotated by humans for open-vocabulary emotions. Following a similar annotation process, OV-MERD+ expands the dataset to a total of 532 samples. In this study, we employ OV-MERD+ as the test set to evaluate the performance of different models on open-vocabulary emotion detection.
MER-UniBench (Lian et al., 2025a).
MER-UniBench is a benchmark that encompasses three representative tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. Specifically, sentiment analysis focuses on binary classification into positive and negative sentiments. Basic emotion recognition restricts the label space to a predefined set of basic emotion categories. In contrast, fine-grained emotion detection removes these constraints, allowing predictions across any category. This benchmark is designed to assess generalized emotion understanding capabilities. Table 6 summarizes the data composition for each task.
Table 6: MER-UniBench. This table presents the data components for three tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. In total, this benchmark includes 12,799 test samples.
Raw Dataset Selected Subset # Samples Label Space
@@ G.2 Testing Datasets
CMU-MOSI (Zadeh et al., 2017) Test 686 positive, negative
CMU-MOSEI (Zadeh et al., 2018) Test 4,659 positive, negative
CH-SIMS (Yu et al., 2020) Test 457 positive, negative
CH-SIMS v2 (Liu et al., 2022) Test 1,034 positive, negative
@@
MER2023 (Lian et al., 2023a) MER-MULTI 411 worry, happy, neutral, angry, surprised, sad
MER2024 (Lian et al., 2024a) MER-SEMI 1,169 worry, happy, neutral, angry, surprised, sad
IEMOCAP (Busso et al., 2008) Session5 1,241 anger, happiness, sadness, neutral
MELD (Poria et al., 2019a) Test 2,610 anger, joy, sadness, neutral, disgust, fear, surprise
@@
OV-MERD+ (Lian et al., 2025a) All 532 unrestricted label space
G.3 Data Leakage Prevention and Overlap Removal
To prevent data leakage, we perform a strict check to ensure that there is no overlap between the training and testing data. During this process, we identify 200 overlapping samples between MER2025-OV (part of the training set) and OV-MERD+ (part of the test set). Consequently, we remove these overlapping samples from MER2025-OV, resulting in 1,000 samples available for training (see Table 1). Apart from this case, we do not observe any other instances of overlap between the training and testing data.
G.4 Task Clarification
In this paper, we focus on utterance-level emotion recognition, as opposed to dialogue-level emotion recognition. The key difference between these tasks lies in whether conversational history is taken into account. For utterance-level emotion recognition, the input consists of a single video clip centered on one character (Tsai et al., 2019). In contrast, dialogue-level emotion recognition requires leveraging additional dialogue information to perform emotion inference (Poria et al., 2017). This distinction has led to the development of two types of benchmarks: the utterance-level benchmark (Lian et al., 2026b) and the dialogue-level benchmark (Poria et al., 2019b).
In MER-UniBench, although certain datasets (e.g., MELD and IEMOCAP) contain dialogue information, others (e.g., MER2023 and CH-SIMS) do not. Furthermore, even when dialogue context is available in datasets like MELD and IEMOCAP, its inclusion is not strictly necessary and depends on the specific experimental setup (Tsai et al., 2019; Poria et al., 2017). Consequently, MER-UniBench is designed as an utterance-level benchmark, and this work adopts the same setting to ensure fair comparison with prior studies (Lian et al., 2025a). Dialogue-level emotion recognition is beyond the scope of this paper and will be explored in our future work.